OpenStudy (anonymous):

a meat company cuts slices of peperoni with a thickness of 3/8 inch if the stick of peperoni is 12 inches long how many slices can be made from 1 stick? *please show work and answer*

563blackghost (563blackghost):

We would turn 12 into a fraction soooo....\[\frac{ 12 }{ 1 }\div \frac{ 3 }{ 8 }=?\] So this is our equation but we need to have similar denominatiors so what would we do?

563blackghost (563blackghost):

If you are looking for just the answer and an explanation for the answer that is against the rules....openstudiers must "help" the studier with their problem and not give away the answer and an explanation to the answer....its against the rules...so plz cooperate

OpenStudy (anonymous):

What blackghost wrote is exactly correct, just divide 12 by 3/8 or .375.

563blackghost (563blackghost):

That is another way we can do it without fractions.....\[12 \div .375=?\]

563blackghost (563blackghost):

Well we would multiply 8 to the fraction 12 and get similar denominators....\[\frac{ 12 }{ 1 }\times \frac{ 8 }{ 8 }=\frac{ 96 }{ 8 }\] now that we have similar denominators we can now divide and solve the problem... That is all I can help with since you are not cooperating...

OpenStudy (anonymous):

@563blackghost i was eating dinner sorry